A balloon is rising with a constant velocity of 10m/s. An object is released freely from that balloon and hits to the ground after 20 seconds. Ignoring the air resistance, find the position of the balloon:
1
Expert's answer
2017-10-25T15:35:07-0400
The equation of motion is
y=h+vt-(gt^2)/2.
The position of the balloon is h.
An object hits to the ground at y=0.
h+vt-(gt^2)/2=0
h=(gt^2)/2-vt=((10) (20)^2)/2-(10)20=1800 m.